Circular No. 7302

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1999 V1 C. W. Hergenrother, Lunar and Planetary Laboratory, reports the dis-

covery of another comet by the Catalina Sky Survey.

1999 UT α_{2000} δ_{2000} m_{1} Observer
Nov. 5.43943 9^{h}37^{m}11.^{s}47 +11^{o}32′03′′.7 17.6 Hergenrother
5.44850 9 37 12.27 +11 31 56.6 17.8  _{”}
5.45751 9 37 12.55 +11 31 52.2 17.5  _{”}
5.46704 9 37 13.18 +11 31 45.8 18.0  _{”}
6.44715 9 38 09.20 +11 22 20.9 18.0  _{”}
6.45209 9 38 09.55 +11 22 18.0 17.6  _{”}
6.47319 9 38 10.79 +11 22 05.8  Shelus
6.48359 9 38 10.98 +11 21 59.7  _{”}
6.49391 9 38 11.62 +11 21 52.8  _{”}

C. W. Hergenrother (Catalina Sky Survey). 0.41-m Schmidt + CCD. Apparent faint tail on single frames. On a 240-s coadded exposure on Nov. 5 there showed a 13′′ coma and a 60′′ tail in p.a. 300^{o}.

P. J. Shelus (McDonald Observatory). 0.76-m reflector with prime focus corrector + LF1 CCD. Fuzzy head and faint tail roughly in p.a. 320^{o}.

COMET P/1999 U3 (LINEAR) The elliptical elements below are from MPEC 1999-V17. That the

comet is of short period has also been pointed out by S. Nakano.

T = 1999 Nov. 5.941 TT  ω = 102.^{o}660
e = 0.65300  Ω = 309.650 }2000.0
q = 1.87220 AU  i = 20.416
a = 5.39542 AU n^{o} = 0.078644 P = 12.53 years

COMET 63P/WILD 1 Nakano also reports observations by T. Kojima, Chiyoda, of this 13-

year-period comet, missed at its 1986 return. These observations confirm a single-night detection at mag 22.4 by Hergenrother (1.5-m Catalina reflector) on Feb. 14. The prediction on MPC 27082 requires correction by ΔT = −0.35 day. Further details are on MPEC 1999-V18. Kojima (0.25-m f/6.3 reflector) reported the comet at m_{1} = 16.5 and as diffuse without a tail on Oct. 24, at m_{1} = 15.9 and diffuse with condensation and a coma diameter of 30′′ on Nov. 4.

1999 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Brian G. Marsden